OpenStudy (anonymous):

Express the required measurement in exact form. If the length of an arc is 12 PI inches and the radius of the circle is 10 inches, what is the measure of the arc? = ? degrees

OpenStudy (anonymous):

circumference = 2*pi*r = 20 pi arclength = 12 pi 12/20 = 3/5 * 360 degrees = 216 degrees

OpenStudy (anonymous):

thanks Can You help me with the next problem

OpenStudy (anonymous):

who knows, try me :)

OpenStudy (anonymous):

lol ok